a wave has an angular frequency of 110 rad/sec and a wave lenth of 1.8m calculate the speed of wave?
 

Grade:12

1 Answers

Kiran Allada
8 years ago
GIven:-
Angular Frequency (ω)P { margin-bottom: 0.08in; })=110rad/sec=110rad/sec
wave length(λ)=1.8m
Frequency(f)=ω/2Π=110/2Π=17.50
speed of the wave(v)=λf=1.8*17.50=31.51m/secP { margin-bottom: 0.08in; }

 

P { margin-bottom: 0.08in; }